OpenStudy (anonymous):

Write the standard equation for the circle that passes through the points: (1, 7) (8, 6) (7, -1) x^2+y^2 ____ x ____ y =0

OpenStudy (baru):

\[x^2 + y^2+ Ax +By +C=0\] plug in those points and find A,B,C

OpenStudy (baru):

or if you prefer, use \[(x-h)^2 + (y-k)^2 = r^2\] and find h,k,r by plugging in those points

OpenStudy (anonymous):

what goes in the spots?

OpenStudy (anonymous):

@baru

OpenStudy (anonymous):

(x+?)+(y-?)=? what numbers would I use from the given ones?

OpenStudy (baru):

you have to substitute for x and y for example (1,7) is a point on the circle so substitute x=1 and y=7 and write down what you get. do this for each of the three points

OpenStudy (anonymous):

(1-7)^2=36 (8-6)^2=4 (7--1)^2=64

OpenStudy (baru):

\[x^2 + y^2+ Ax +By +C=0\] substitute x=1 and y=7 we get \[1 + 49 +A + 7B +C=0\] which simplifies to \[A +7B +C +50 =0\]

OpenStudy (baru):

you have to do the same procedure with the other two points, and get two more equations in A,B,C

OpenStudy (anonymous):

okay